A 66-foot board is to be cut into two pieces whose lengths are in the ratio of 4 to 7. Find the lengths of the two pieces.
Answer by jim_thompson5910(35256)

The ratio is 4:7 so that means the two pieces are 4x and 7x ft long. They add to 66. So,


4x + 7x = 66


11x = 66


x = 66/11


x = 6


--------------


4x = 4*6 = 24


7x = 7*6 = 42


--------------


The lengths of the two pieces are 24 ft and 42 ft